Operations Report – April 17th

Crew 280 Operations Report 16-04-2023
SOL: 1
Name of person filing report: Cesca Cufí – Crew Engineer (assisted by Neus Sabaté – Mission Specialist)
Non-nominal systems: robotic observatory
Notes on non-nominal systems: Replacement parts are scheduled to be installed during crew 261 rotation.
Spirit rover used:
Hours: 216.6
Beginning charge: 100%
Ending charge: 93%
Currently charging: no
Opportunity rover used:
Hours: 120.7
Beginning charge: 100%
Ending charge: 100%. We confirm the rover display was showing 100% battery charge after the EVA.
Currently charging: no
Curiosity rover used:
Hours: 229.7
Beginning charge: 100%
Ending charge: 99%
Currently charging: no
Perseverance rover used:
Hours: 261.1
Beginning charge: 100%
Ending charge: 96%
Currently charging: no
General notes on rovers: non
Summary of Hab operations:
Water use: can not be estimated as water content has increased
Water (static tank): 229 gallons
Static tank pipe heater (on or off): off
Static tank heater (On or off): off
Toilet tank emptied: no
Summary of internet: used from 18:00 to 21:00 for personal communication and reports.
Summary of suits and radios:
Suits: space suit number 2 found with the collar mounted upside down and with only one metal bridle (see attached picture “spacesuit_2”). It has been repaired with a spare metallic bridle.
Radios: short fail in communication channel 1 while in EVA to Marble Ritual on radio 2 and channel 2 on radio 5.
Radio 2: the COMMS EVA responsible, Mariona, was not able to hear the COMMS Station responsible on channel 1 for a couple of minutes. The issues appeared when the crew member Mariona was far away from the station but still on sight from the station (no terrain obstruction possible). The operation of channel 1 was switched to crew member Carla, the backup responsible. She maintained communication with the Station and reported the issue. The COMMS on channel 1 with radio 2 were recovered without specific troubleshooting. No troubleshooting was tried. The second EVA didn’t report any issues operating on the same channel around the same location.
Radio 5: the COMMS on channel 2 had some issues. The crew member was pressing the PTT button but the voice was not heard after some seconds. That is, there was a delay on the activation of the voice message (greater than the common delay). This issue appears when the crew member operating the radio was already back to the station, next to the Hub, and was getting off the rover. The crew member being close to the station and ready to enter the Hub, no troubleshooting was tried.
We will troubleshoot the future radio issues during the EVA on-site.
Both radios were working fine during pre-EVA check and most of the EVA time. Both radios have been tested after EVA and they work nominally.
We highly suspect a misconnection on the headsets plug, maybe due to the movement of the crew members.
All radio and spacesuit batteries checked before EVA: ok.
Radio batteries charged after EVA.
Spacesuit batteries plugged.
Summary of GreenHab operations:
Water use: 13 gallons
Heater: off
Supplemental light: off
Harvest: 1g (thyme, sage), 114 g microgreens, 56 g cherry tomatoes
Summary of Science Dome operations:
Transport of science material to the dome.
Dual split: off
Summary of RAM operations: none
Summary of any observatory issues: none
Robotic observatory: Not working. Replacement parts are scheduled to be installed during crew 261 rotation.
Summary of health and safety issues:
Cesca, Ari and Núria were back at the station at 11.30h this morning. Crew Engineer Cesca Cufí has been checked in Moab Regional Hospital for abdominal pain. After some exams: appendicitis discarded, it may be an intestinal infection. Cesca was given some antibiotics at the hospital last night and prescribed some antibiotics and pain killers for the next 5 days. Cesca feels a lot better, still with some pain but functional.

Questions, concerns and requests to Mission Support:
Spacesuit 2: to properly repair the spacesuit, we’ve searched for the metallic bridles in the RAM but we couldn’t find any. Do you confirm there are no spare bridles of this king in the station? If any, we will replace the plastic bridle for the metallic one. Thanks.
Update: The plastic bridle has been replaced by the metallic one.

Copyright © The Mars Society. All rights reserved. | Main Site